OpenStudy (anonymous):

Multiply (12u3 / 7v5) * (14v2 / 3u)

OpenStudy (anonymous):

\[\frac{12 u^3}{7 v^5}*\frac{14 v^2}{3 u}=\frac{8 u^2}{v^3} \]